OpenStudy (anonymous):

The volume of water in a bowl is given by V = 1/3(pi)(h)^2(60 - h) cm^3, where h is the depth of the water in centimeters. If the depth is increasing at the rate of 4 cm/sec when the water is 10 cm deep, how fast is the volume increasing at the instant?

OpenStudy (lastdaywork):

V = (1/3)pi(60h^2 - h^3) dh/dt(at h=10) = 4 Find dV/dt (at h=10)

OpenStudy (lastdaywork):

Can you solve the question now ??

OpenStudy (anonymous):

Let me see. One second

OpenStudy (anonymous):

5000pi/3 right?

OpenStudy (lastdaywork):

mine was 1200 pi :(

OpenStudy (anonymous):

Oh, how did you get it? I must have simplified! :D

OpenStudy (anonymous):

Well. where do we go after finding 1200pi?

OpenStudy (lastdaywork):

Would you mind giving me your solution ?

OpenStudy (anonymous):

I'm on a study guide. I thought 5000pi/3 (the wrong one) was the final answer but I think we might need to do implicit diffs, right?

OpenStudy (lastdaywork):

Post your solution for this question..I'll correct it :)

OpenStudy (anonymous):

That's the only solution I had, 5000pi/3 :) This is a real confusin section :P

OpenStudy (lastdaywork):

I mean; show me your steps..how you reached the answer..

OpenStudy (anonymous):

Well for the wrong one i just simplified (1/3)(pi)(60(10)^2-10^3)

OpenStudy (lastdaywork):

You need to find dV/dt at h = 10 So differentiate first..then substitute h

OpenStudy (anonymous):

ok let me see if i can

OpenStudy (anonymous):

so 20pi^2-pi^2(h)/3 20pi^2-pi^2(10)/3 50pi^2/3

OpenStudy (lastdaywork):

pi^2 ??

OpenStudy (anonymous):

ya, im a little confused. my other answer still had p^2 in it? am i off?

OpenStudy (lastdaywork):

pi is a constant..

OpenStudy (lastdaywork):

so are you differentiating a constant ??

OpenStudy (anonymous):

i might be. I'm kinda lost

OpenStudy (lastdaywork):

V = (1/3)pi(60h^2 - h^3) dv/dt = (1/3)pi(120h - 3h^2) * dh/dt dv/dt = 1200 pi Does that makes sense ??

OpenStudy (anonymous):

ok a couple questions. where does th120 and 3 come from?

OpenStudy (lastdaywork):

Do you know differentiation ??

OpenStudy (anonymous):

finding derivative right?

OpenStudy (lastdaywork):

Yea..what's the derivative of x^n ??

OpenStudy (anonymous):

nx^n-1

OpenStudy (lastdaywork):

Now, do you know Chain Rule ??

OpenStudy (anonymous):

i think. Should i use it in this problem?

OpenStudy (lastdaywork):

Whats the derivative of x^n with respect to t ?

OpenStudy (anonymous):

I'm not sure

OpenStudy (lastdaywork):

d(x^n)/dt = [d(x^n)/dx] * [dx/dt] = nx^(n-1) * dx/dt Chain Rule ^^ :)

OpenStudy (anonymous):

ok. so am i going to apply it to this one? I'm not all that sure about where to go after 1200pi

OpenStudy (lastdaywork):

1200*pi is the rate at which volume is increasing at h=10

OpenStudy (anonymous):

so, the final is 1200pi?

OpenStudy (lastdaywork):

yep..

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

I don't think because it needs to be in cm^3/sec

OpenStudy (lastdaywork):

yea..it is in cm^3/sec

OpenStudy (anonymous):

that's still not matching up. Let me write the equation out for you. holdon

OpenStudy (anonymous):

1/3(pi)(h)^2(60 - h) \[1/3pih^2(60 -h)\]

OpenStudy (lastdaywork):

"... V = (1/3)pi(60h^2 - h^3) dv/dt = (1/3)pi(120h - 3h^2) * dh/dt dv/dt = 1200 pi ..." IDK what else to explain :(

OpenStudy (anonymous):

Ok. 1 second let me see if i can verify it. @sourwing

OpenStudy (lastdaywork):

@agent0smith Can you help here please..

OpenStudy (anonymous):

thanks for calling someone. sour might be away

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles \[V=\frac{1}{3}\pi h^2(60-h)\] You're given \(\dfrac{dh}{dt}=4\), and you're to find \(\dfrac{dV}{dt}\) when \(h=10\). Differentiate both sides with respect to \(t\) (I use the product rule, but nothing is stopping you from expanding the polynomial and just using the power rule term-by-term): \[\frac{dV}{dt}=\frac{\pi}{3}\left[2h(60-h)\frac{dh}{dt}+h^2(-1)\frac{dh}{dt}\right]\] Plugging in the given values, you have \[\frac{dV}{dt}=\frac{\pi}{3}\left[2(10)(60-10)(4)+10^2(-1)(4)\right]=\frac{\pi}{3}(4000-400)=\color{red}{1200\pi}\frac{\text{cm}^3}{\text{s}}\] \(\color{blue}{\text{End of Quote}}\)